Fermi National Laboratory


Accelerator Update

Friday May 21
The day shift began with the Tevatron (TeV) recovering from a quench, and with MiniBooNE taking beam.

Operations prepared the Collider Detector at Fermilab (CDF) and the D-zero (D0) collision halls for access at 8:50 AM.

The Antiproton Source (Pbar) began stacking at 9:04 AM.

TeV experts began a study period around noon.

Operations reported at 1:07 PM that a Linac quadrupole power supply was going into and out of alarm. An expert switched out some control cards, but the supply was still showing trouble at 1:44 PM.

A TeV RF station (TRF2) tripped off at 1:07 PM.

A TeV RF station (TRF3) tripped off at 2:05 PM.

Operations established beam to the Meson MTest experiment 2:42 PM.

A CDF experimenter reported that their chilled water system tripped off at 3:57 PM. Operations delayed putting the TeV into shot setup while experimenters investigated the problem.

At 4:09 PM, a Booster RF station (BRF11) and a power supply (GMPS) tripped off on high water temperature. Operators investigated.

Booster experts needed to access their tunnel so operators prepared the area for access at 4:31 PM. At 5:16 PM, Booster experts reported finding a bad klixon. They bypassed the klixon.

TeV experts completed their studies and Operations put the TeV through a dry squeeze at 6:40 PM.

At 6:53 PM, Operations reported problems getting beam out of Linac and to the Booster. Operators called in a Linac expert.

Between 8 and 9 PM repairs and investigations were delayed due to tornado warning evacuations.

Both CDF and D0 experimenters requested access to their collision hall at 9:16 PM.

MTest finished using the beam for the night at 9:37 PM.

At 10:02 PM, a lightning strike interfered with the TeV ramp and tripped off some TeV RF stations, but there was no quench.

D0 experimenters came out of access at 10:21 PM, but CDF experimenters said they needed two more hours.

Booster experts repaired the Linac 400 MeV chopper and various other components. Operations established beam to Pbar at 11:40 PM.

Saturday May 22
The midnight shift began with CDF in access, Pbar stacking, and with MiniBooNE and MCenter taking beam.

CDF came out of access at 1:09 AM and Operations put the TeV through a dry squeeze.

Operations had started to put the TeV through a wet squeeze when an RF problem aborted the beam at 1:53 AM. Operators investigated and discovered that two RF stations (TRF7 & 8) weren't putting out RF. An expert conducted a full gap envelope reset and both stations resumed normal operations.

Operations put the TeV into shot setup at 3:23 AM. All Switchyard beam (SY) gets put on hold during shot setups.

At 4:29 AM, the Pbar beam suffered an emittance blowup. Experts had the situation fixed thirty minutes later.

At 4:31 AM, the Main Injector (MI) suffered from a coalescing problem. At 5:16 AM, operators discovered a problem with the dampers. They turned off the dampers and turned on the Booster Fuzzer (which is a damper), but coalescing still looked bad. An expert came in to investigate.

The Run Coordinator decided stacking should resume, but when Operations moved Pbar from the shot-lattice to the stacking-lattice the stack dropped from 117mA to 38mA due to an emittance instability that occurred during the switch.

Pbar resumed stacking at 7:54 AM.

Operations established beam to MCenter at 9:12 AM and to MTest at 10:52 AM.

Operations put the TeV through a wet squeeze from 12:30 PM to 12:59 PM.

D0 experimenters requested access to their collision hall. Operators had the area prepared for access by 1:32 PM.

At 1:58 PM, Operations reported that a Pbar Debuncher extraction kicker (D:EKIK) tripped off on an oil flow problem. Operators reset the device, but it tripped off again. Operators called in an expert.

D0 came out of access at 3:33 PM.

A Pbar expert replaced a faulty power supply module on D:EKIK and stacking resumed at 3:52 PM.

Operations put the TeV into shot setup at 4:45 PM and at 6:17 PM established store 3528 with an initial luminosity of approximately 47.46E30. (CDF had a problem with measuring their luminosity.)

Pbar resumed stacking at 6:35 PM. MCenter and MTest resumed taking beam.

A MCenter expert began tuning the beam line at 9:47 PM. She finished about two hours later.

Sunday May 23
The midnight shift began with Operation monitoring store 3528, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 3:06 AM, CDF requested Operations to fix a colliding beam problem. Operators both tuned and moved a collimator.

A Recycler expert began a study period from 7:37 AM to 10:54 AM.

Earlier, a Booster RF problem interrupted beam. At 8:31 AM, operators discovered a problem with BRF1. They bypassed the station and called an expert.

An MI expert began a four and one half hour barrier bucket study at 11:04 AM.

At 6:08 PM, a Linac RF station, which had been having problems during the shift, tripped off. It tripped off two more times in ten minutes.

Operations terminated store 3528 at 6:16 PM and put the TeV into shot setup at 7 PM.

Operations reported at 7:17 PM that a Pbar emittance blowup was delaying shot setup.

Operations established store 3530 at 10 PM with an initial luminosity of 66.44E30. Resumed sending beam to SY.

At 11:11 PM, Operation reported they had a problem getting Pbar to resume stacking due to a cable interlock problem. Operations bypassed the device after all resets failed.

Monday May 24
The midnight shift began with Operations monitoring store 3530, with Pbar stacking, with MiniBooNE and MCenter taking beam.

At 1:11 AM, a Linac RF station (KRF5) tripped off the first of four times over ninety minutes.

At 3:33 AM, a Linac RF station (KRF2) tripped off and wouldn't reset. Operators investigated and discovered a water flow problem. They convinced the flowmeter that it would be better off to work normally and eventually, at 3:57 AM, it agreed.

MCenter experimenters decided they needed an extended access so Operations turned off their beam at 6:05 AM.

A Recycler expert began a study period at 6:20 Am.

At 7:24 AM, Operations reported that a Linac quadrupole power supply (L:QPS111) had been tripping regularly. Operators tried to improve its internal airflow, which helped a little.

Plans for Today
The plan for today allows Pbar experts to conduct a Debuncher study this morning that will disrupt stacking and for MI studies.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/24/2004   email Fermilab